# Analysis and Synthesis Denoisers

### v2

@@ 3.2 Synthesis denoiser structure
$\begin{array}[]{l}\text{for }\ell=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}z_{\ell+1}=\operatorname{prox}_{\zeta g_{\lambda}}\big{(}z_{\ell}-\zeta D^{*}(Dz_{\ell}-v)\big{)},\end{array}\right.\end{array}$ (3.16)
where $z_{0}\in\mathbb{R}^{S}$, and $\zeta>0$ is a step-size. Then the following convergence result can be deduced from [17, Thm. 3.4].
@@
Assume that there exists $\underline{\zeta}>0$ such that ${\underline{\zeta}<\zeta<(2-\underline{\zeta})\|D\|^{-2}_{S}}$. Then the sequence ${(z_{\ell})}_{\ell\in\mathbb{N}}$ converges to $z_{D}^{\dagger}$, and $x^{\dagger}_{D}=Dz_{D}^{\dagger}$ is a solution to (3.1).
algorithm (3.16) can be used as sub-iterations when included within the global FB algorithm (1.5), to approximate the computation of $\operatorname{prox}_{D\rhd g_{\lambda}}$. Similarly to Section 2, in this section we investigate the behavior of (1.5) with the SD, when only one iteration of algorithm (3.16) is computed. To this aim, we define an unfolded iterative scheme approximating the proximal operator of $p=D\rhd g_{\lambda}$, i.e., an SD as defined in (1.7).
@@
[…]
@@
    $(\forall z\in\mathbb{R}^{S})\quad T^{\mathcal{S}}_{\lambda,v}(z)=\operatorname{prox}_{\zeta g_{\lambda}}\Big{(}\big{(}\operatorname{Id}-\zeta D^{*}D\big{)}z+\zeta{D}^{*}v\Big{)}.$
@@
    A direct consequence of Theorem 3.4 is that, for any $z_{0}\in\mathbb{R}^{S}$, if there exists $\underline{\zeta}>0$ such that $\underline{\zeta}\leq\zeta\leq(2-\underline{\zeta})\|D\|_{S}^{-2}$, then $\lim_{L\to+\infty}G^{\mathcal{S}}_{L,\lambda,v}(z_{0})=x_{D}^{\dagger}$, where $x_{D}^{\dagger}$ is defined in (3.1).
@@ 3.3 FB-PnP algorithm with approximated SD
$\begin{array}[]{l}x_{0}\in\mathbb{R}^{N},z_{0}\in\mathbb{R}^{S},\\ \text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}{v}_{k}=x_{k}-\tau A^{*}(Ax_{k}-y),\\ z_{k+1}={G}^{\mathcal{S}}_{L,\tau\lambda,v_{k}}(z_{k}),\\ x_{k+1}=Dz_{k+1},\end{array}\right.\end{array}$ (3.20)
@@
Let ${(x_{k})}_{k\in\mathbb{N}}$ be generated by algorithm (3.20). Assume that there exist $\underline{\tau}>0$ and $\underline{\zeta}>0$ such that $\underline{\tau}\leq\tau\leq(2-\underline{\tau})\|A\|_{S}^{-2}$ and $\underline{\zeta}<\zeta<(2-\underline{\zeta})\|D\|_{S}^{-2}$. When $L\to\infty$, i.e. i.e. when we use an infinite number of iterations in Model 3.5, then ${(x_{k})}_{k\in\mathbb{N}}$ converges to a solution to problem (3.6).
Theorem 3.7 only holds when $L\to\infty$, i.e. when $G^{\mathcal{S}}$ has a very large number of sub-iterations (i.e., layers). As for the AD case, this is often intractable, and only a fixed (low) number of iterations are considered in practice. In the remainder of this section, we focus on the case when only $L=1$ sub-iteration is computed in Model 3.5, and investigate the behavior of ${(x_{k})}_{k\in\mathbb{N}}$ in this context, as well as the nature of the provided solution. In this particular case, at iteration $k\in G$, the update of $z_{k}$ in algorithm (3.20) boils down to
$z_{k+1}={G}^{\mathcal{S}}_{1,\tau\lambda,v_{k}}(z_{k})=T^{\mathcal{S}}_{\tau\lambda,v_{k}}(z_{k}),$ (3.21)
@@
Let ${(z_{k})}_{k\in\mathbb{N}}$ and ${(x_{k})}_{k\in\mathbb{N}}$ be sequences generated by algorithm (3.20) with $L=1$. Assume that there exists $\varepsilon>0$ such that $\varepsilon\leq\tau\zeta\leq(2-\varepsilon)\|AD\|_{S}^{-2}$, then
@@
$x_{k+1}$ $=Dz_{k+1}.$ (3.25)
Combining (3.22) and (3.23), and then using (3.25), we obtain
@@
$\begin{array}[]{l}\text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}\widetilde{z}_{k}=z_{k}-\tau\zeta D^{*}A^{*}(ADz_{k}-y),\\ z_{k+1}=\operatorname{prox}_{\tau\zeta g_{\lambda}}(\widetilde{z}_{k}),\\ x_{k+1}=Dz_{k+1}.\end{array}\right.\end{array}$ (3.26)
algorithm (3.26) corresponds to the FB iterations given in (3.8) with step-size $\gamma=\tau\zeta$. Then, according to [17, Thm. 3.4], ${(z_{k})}_{k\in\mathbb{N}}$ converges to a solution $\widehat{z}_{S}$ to (3.7) if there exists $\varepsilon>0$ such that $\varepsilon\leq\tau\zeta\leq(2-\varepsilon)\|AD\|_{S}^{2}$. Further, according to Theorem 3.2(i), $\widehat{x}_{S}=D\widehat{z}_{S}$ is a solution to (3.6).
To summarize, in Theorem 3.8, as in the case of AD, we show that we can recover the solution of the full problem (3.6) using only $L=1$ sub-iterations, provided that we use a restarting scheme.
@@ 4 Approximated AD and SD for smooth minimization
where $\widetilde{f}\colon\mathbb{R}^{\widetilde{N}}\to]-\infty,+\infty]$ is assumed to be convex and $\widetilde{\beta}$-Lipschitz differentiable, for $\widetilde{\beta}>0$, and $\widetilde{g}_{\lambda}\colon\mathbb{R}^{\widetilde{N}}\to]-\infty,+\infty]$ is a convex, proper and lower-semicontinuous function. Then, the synthesis problem is a particular case of (4.1) where
$\begin{cases}\begin{array}[]{l@{}l@{}l}\widetilde{f}\colon&\mathbb{R}^{S}\to]-\infty,+\infty]\colon&z\mapsto\|ADz-y\|^{2}\\ \widetilde{g}_{\lambda}\colon&\mathbb{R}^{S}\to]-\infty,+\infty]\colon&z\mapsto g_{\lambda}(z).\end{array}\end{cases}$ (4.2)
@@
$\begin{cases}\begin{array}[]{l@{}l@{}l}\widetilde{f}\colon&\mathbb{R}^{N}\to]-\infty,+\infty]\colon&x\mapsto\|Ax-y\|^{2}\\ \widetilde{g}_{\lambda}\colon&\mathbb{R}^{N}\to]-\infty,+\infty]\colon&x\mapsto g_{\lambda}(\Gamma x).\end{array}\end{cases}$ (4.3)
@@
We would emphasize that, according to Theorem 3.2, if $\widehat{z}$ is a solution to (4.1)-(4.2), then $D\widehat{z}$ is a solution to (3.6).
@@
find $\text{find }x^{\ddagger}=\underset{x\in\mathbb{R}^{\widetilde{N}}}{\operatorname{argmin}}\;\widetilde{f}(x)+{{}^{\mu}\widetilde{g}_{\lambda}}(x),$ (4.4)
@@
$\begin{array}[]{l}x_{0}\in\mathbb{R}^{N},u_{0}\in\mathbb{R}^{N},\\ \text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}x_{k+1}=x_{k}-\tau\nabla_{x}F(x_{k},u_{k}),\\ u_{k+1}=\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x_{k+1}),\end{array}\right.\end{array}$ (4.5)
where, for every $(x,u)\in\mathbb{R}^{\widetilde{N}}\times\mathbb{R}^{\widetilde{N}}$, $F(x,u)=\widetilde{f}(x)+\widetilde{g}_{\lambda}(u)+\frac{\mu}{2}\|x-u\|^{2}$, and
@@
is Lipschitz-differentiable, with constant $\widetilde{\beta}+\mu$. The sequence ${(x_{k})}_{k\in\mathbb{N}}$ generated by algorithm (4.5) is then ensured to converge to the solution $x^{\ddagger}$ to (4.4), if there exists $\underline{\zeta}>0$ such that $\underline{\zeta}\leq\tau\leq(2-\underline{\zeta}){(\widetilde{\beta}+\mu)}^{-1}$.
In practice, as the proximity operator of $\widetilde{g}_{\frac{\lambda}{\mu}}$ often does not have a closed form, the computational complexity of this algorithm is too high. However, the solution of the proximity operator can be computed using iterative schemes such as the FB or the dual-FB algorithms, as described in Sections 2 and 3. Using a bilevel framework similar to the one proposed by [19], we show that one can use inexact proximity operator computation powered by iterative proximal algorithms with warm restarts to build an algorithm that converges to a solution to (4.4). In this context, we then focus on the following algorithm
$\begin{array}[]{l}x_{0}\in\mathbb{R}^{N},u_{0}\in\mathbb{R}^{N},\\ \text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}x_{k+1}=x_{k}-\tau\nabla_{x}F(x_{k},u_{k}),\\ u_{k+1}=\widetilde{G}_{L,\lambda\mu^{-1}}(x_{k+1},u_{k}),\end{array}\right.\end{array}$ (4.7)
where, for every $k\in\mathbb{N}$, $\widetilde{G}_{L,\lambda\mu^{-1}}\colon\mathbb{R}^{\widetilde{N}}\times\mathbb{R}^{\widetilde{N}}\to\mathbb{R}^{\widetilde{N}}$ is an operator that aims to approximate $\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x_{k+1})$ by computing $L$ sub-iterations of some iterative algorithm. The first input corresponds to the output $x_{k+1}$ of the gradient descent step, i.e. the point at which the proximity step should be computed. The second input corresponds to the output of $\widetilde{G}_{L,\lambda\mu^{-1}}$ from the previous iteration $k-1$, which is aimed to be used for warm-restart, i.e. initializing the sub-iterations to compute the approximation of $\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x_{k+1})$. Unlike in previous sections, in this section, this operator is not necessarily based on FB or dual-FB iterations. However, we will assume that it satisfies some sufficient decrease property (see Theorem 4.3 Condition (i)).
@@
$\text{find }x^{\ddagger}=\underset{x\in\mathbb{R}^{\widetilde{N}}}{\operatorname{argmin}}\;\widetilde{f}(x)+\widetilde{g}_{\lambda}(u^{\ddagger}_{x})+\frac{\mu}{2}\|x-u^{\ddagger}_{x}\|^{2}\quad\text{such that}\quad u^{\ddagger}_{x}=\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x),$ (4.8)
where $\mu>0$. Hence, the sequence ${(x_{k})}_{k\in\mathbb{N}}$ generated by (4.5) is ensured to converge to a solution to (4.8) if there exists $\underline{\tau}>0$ such that $\underline{\tau}\leq\tau\leq(2-\underline{\tau}){(\widetilde{\beta}+\mu)}^{-1}$.
@@
    $\phi\in\mathbb{R}_{*}\mapsto p(\phi)=(8\alpha_{L}^{2}(1-2\alpha_{L}^{2}))\phi^{2}+2(\widetilde{\beta}+\mu)(1-2\alpha_{L}^{2})\phi-\mu^{2}.$ (4.10)
@@
Proof. Let $k\in\mathbb{N}$, and let us define, for every $x\in\mathbb{R}^{\widetilde{N}}$, $h(x)=F(x,u^{\ddagger}_{x})$, where $u^{\ddagger}_{x}\in\mathbb{R}^{\widetilde{N}}$ is defined in (4.8). According to (4.6), since $\nabla_{x}F(\cdot,u^{\ddagger}_{x})$ is $(\widetilde{\beta}+\mu)$-Lipschitz, then $h$ is Lipschitz-differentiable, with constant $\widetilde{\beta}+\mu>0$. Thus we can apply the descent lemma from [6, Prop. A.24] to obtain
@@
$\|u_{k+1}-u^{\ddagger}_{x_{k+1}}\|\leq\alpha_{L}\|u_{k}-u^{\ddagger}_{x_{k}}\|+\alpha_{L}\tau\|\nabla_{x}F(x_{k},u_{k})\|,$
and by applying the Jensen’s inequality we obtain
@@
Hence, by summing (4.19) over $k\in\{0,\ldots,K-1\}$, for $K>0$, we obtain
$\sum_{k=0}^{K}\frac{\tau}{2}\|\nabla h(x_{k})\|^{2}\leq\mathcal{L}_{0}-\mathcal{L}_{K}<+\infty.$
Thus ${(\|\nabla h(x_{k})\|^{2})}_{k\in\mathbb{N}}$ converges to 00. Since $h$ is convex with continuous gradient, the gradient vanishes at a global minimum, and ${(x_{k})}_{k\in\mathbb{N}}$ converges to a solution to (4.8).
@@
    Condition (i) in Theorem 4.3 means that the operator $\widetilde{G}_{L,\lambda\mu^{-1}}$ in (4.7) must be chosen to satisfy some sufficient decrease property in the sense that, for every $k\in\mathbb{N}$, $u_{k+1}=\widetilde{G}_{L,\lambda\mu^{-1}}(x_{k+1},u_{k})$ should be closer to $\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x_{k+1})$ than $u_{k}$. In other words, this means that $\widetilde{G}_{L,\lambda\mu^{-1}}$ decreases with respect to its second variable (i.e., the initialisation of that operator).
    In practice, if we choose a monotone algorithm such as FB or dual-FB, this condition can be satisfied for an $L$ sufficiently large. However, it can be difficult to determine the value of $L$ ensuring $\alpha_{L}<1/\sqrt{2}$.
@@
    For $(x,u)\in(\mathbb{R}^{N})^{2}$, we define $\bar{G}_{\lambda}(x,u)=\frac{1}{2}\|x-u\|+\widetilde{g}_{\lambda}(u)$. If $L$ iterations of the algorithm that estimates $u^{\ddagger}_{x}=\operatorname{prox}_{\widetilde{g}_{\lambda}}(x)=\min_{u}\bar{G}_{\lambda}(x,u)$, warm started with a value $u_{\ell}$ produces an estimate $u_{\ell+L}$ such that
@@
    where $C>0$ is a constant that does not depend on $x$, then, by $\lambda^{-1}$-strong convexity of $\bar{G}_{\lambda}(x,\cdot)$ we have
    $\|u_{\ell+L}-u^{\ddagger}_{x}\|_{2}^{2}\leq\frac{2\lambda C}{L}\|u_{\ell}-u^{\ddagger}_{x}\|^{2}.$
    In this case, $L$ should be chosen larger than $4\lambda C$ to satisfy assumption (4.9). Note that this is typically the case when $g$ is the SD formulation and we use FB as a inner solver [5].
@@
    When $\alpha_{L}$ is unknown, the root of the polynomial (4.10) cannot be computed. In this case, one can choose $\phi=1$ in the Lyapunov function (4.18), and ${0<\tau<\min\Big{\{}\frac{2(1-2\alpha_{L}^{2})}{\mu^{2}},\frac{1}{\widetilde{\beta}+\mu+4\alpha_{L}^{2}}\Big{\}}}$. Since $20<\alpha_{L}^{2}<1/$ 2, if $\tau$ is chosen to satisfy ${0<\tau<\min\big{\{}2/\mu^{2},1/(\widetilde{\beta}+\mu)\big{\}}}$, then ${(x_{k})}_{k\in\mathbb{N}}$ generated by algorithm (4.7) converges to a solution to (4.8).
    In practice, since $\mu$ is the smoothing parameter, it is often chosen very small. In this context, we recover a similar condition as for the convergence of (4.5), that is $0<\tau<1/(\mu+\widetilde{\beta})$.

### v3

@@ 3.2 Synthesis denoiser structure
[…]
where $z_{0}\in\mathbb{R}^{S}$, and $\zeta>0$ is a step-size. Then the following convergence result can be deduced from [20, Thm. 3.4].
@@
Assume that $0<\zeta<2\|D\|^{-2}_{S}$}. Then the sequence $(z_{\ell})$$_{\ell\in\mathbb{N}}$ converges to $z_{D}^{\dagger}$, and $x^{\dagger}_{D}=Dz_{D}^{\dagger}$ is a solution to (3.1).
Algorithm (3.16) can be used as sub-iterations when included within the global FB algorithm (1.5), to approximate the computation of $\operatorname{prox}_{D\rhd g_{\lambda}}$. Similarly to Section 2, in this section we investigate the behavior of (1.5) with the SD, when only one iteration of algorithm (3.16) is computed. To this aim, we define an unfolded iterative scheme approximating the proximal operator of $p=D\rhd g_{\lambda}$, i.e., an SD as defined in (1.7). In the remainder, we use the superscript $\mathcal{S}$ to identify operators associated with the synthesis formulation.
@@
$\widetilde{G}^{\mathcal{S}}_{L,\lambda,v}(z_{0})=\underset{\text{$L$ compositions}}{\underbrace{T^{\mathcal{S}}_{\lambda,v}\circ\dots\circ T^{\mathcal{S}}_{\lambda,v}}}(z_{0}),$ (3.17)
@@
    $(\forall z\in\mathbb{R}^{S})\quad T^{\mathcal{S}}_{\lambda,v}(z)=\operatorname{prox}_{\zeta g_{\lambda}}\Big{(}\big{(}\operatorname{Id}-\zeta D^{*}D\big{)}z+\zeta{D}^{*}v\Big{)},$
    where $\operatorname{Id}-\zeta D^{*}D$ is a linear operator and $\zeta{D}^{*}v$ a bias. This reformulation highlights the residual structure of each layer, and makes explicit the operator and bias involved in the synthesis denoiser.
@@
    A direct consequence of Theorem 3.4 is that, for any $z_{0}\in\mathbb{R}^{S}$, if $0<\zeta<2\|D\|_{S}^{-2}$, then $\lim_{L\to+\infty}G^{\mathcal{S}}_{L,\lambda,v}(z_{0})=x_{D}^{\dagger}$, where $x_{D}^{\dagger}$ is defined in (3.1).
@@ 3.3 FB-PnP algorithm with approximated SD
[…]
@@
Let ${(x_{k})}_{k\in\mathbb{N}}$ be generated by algorithm (3.20). Assume that $0<\tau<2\|A\|_{S}^{-2}$ and $0<\zeta<2\|D\|_{S}^{-2}$. When $L\to\infty$ in Model 3.5 (i.e., when problem (3.1) is solved accurately) then ${(x_{k})}_{k\in\mathbb{N}}$ converges to a solution to problem (3.6).
Similarly to Theorem 2.4, Theorem 3.7 only holds when $L\to\infty$, i.e. when $\widetilde{G}^{\mathcal{S}}$ has a very large number of sub-iterations (i.e., layers). As for the AD case, this is intractable, and only a fixed (low) number of iterations are considered in practice. In the remainder of this section, we focus on the case when only $L=1$ sub-iteration is computed in Model 3.5, and investigate the behavior of ${(x_{k})}_{k\in\mathbb{N}}$ in this context, as well as the nature of the provided solution. In this particular case, at iteration $k\in\mathbb{N}$, the update of $z_{k}$ in algorithm (3.20) boils down to
$z_{k+1}=\widetilde{G}^{\mathcal{S}}_{1,\tau\lambda,v_{k}}(z_{k})=T^{\mathcal{S}}_{\tau\lambda,v_{k}}(z_{k}),$ (3.21)
@@
Let ${(z_{k})}_{k\in\mathbb{N}}$ and ${(x_{k})}_{k\in\mathbb{N}}$ be sequences generated by algorithm (3.20) with $L=1$. Assume that $0<\tau\zeta<2\|AD\|_{S}^{-2}$, then
@@
$x_{k+1}$ $=Dz_{k+1},$ (3.25)
where (3.24) is obtained by definition of $T^{\mathcal{S}}_{\tau\lambda,v_{k}}(z_{k})=\operatorname{prox}_{\zeta g_{\tau\lambda}}(\widetilde{z}_{k})$, and noticing that $g_{\lambda\tau}=\tau g_{\lambda}$. Combining (3.22) and (3.23), and then using (3.25), we obtain
@@
[…]
Algorithm (3.26) corresponds to the FB iterations given in (3.8) with step-size $\gamma=\tau\zeta$. Then, according to [20, Thm. 3.4], ${(z_{k})}_{k\in\mathbb{N}}$ converges to a solution $\widehat{z}_{S}$ to (3.7) if $0<\tau\zeta<2\|AD\|_{S}^{2}$. Further, according to Theorem 3.2(i), $\widehat{x}_{S}=D\widehat{z}_{S}$ is a solution to (3.6).
To summarize, in Theorem 3.8, as in the case of AD, we show that we can recover the solution of the full problem (3.6) using only $L=1$ sub-iterations, provided that we use a warm-restarting strategy.
@@ 4 Approximated AD and SD for smooth minimization
[…]
$\begin{cases}\begin{array}[]{l@{}l@{}l}\widetilde{f}\colon&\mathbb{R}^{S}\to(-\infty,+\infty]\colon&z\mapsto\|ADz-y\|^{2}\\ \widetilde{g}_{\lambda}\colon&\mathbb{R}^{S}\to(-\infty,+\infty]\colon&z\mapsto g_{\lambda}(z).\end{array}\end{cases}$ (4.2)
@@
$\begin{cases}\begin{array}[]{l@{}l@{}l}\widetilde{f}\colon&\mathbb{R}^{N}\to(-\infty,+\infty]\colon&x\mapsto\|Ax-y\|^{2}\\ \widetilde{g}_{\lambda}\colon&\mathbb{R}^{N}\to(-\infty,+\infty]\colon&x\mapsto g_{\lambda}(\Gamma x).\end{array}\end{cases}$ (4.3)
@@
We would emphasize that, according to Theorem 3.2, if $\widehat{z}$ is a solution to full synthesis problem (4.1)-(4.2), then $D\widehat{z}$ is a solution to analysis problem with SD (3.6).
@@
$\text{find }x^{\ddagger}=\underset{x\in\mathbb{R}^{\widetilde{N}}}{\operatorname{argmin}}\;\widetilde{f}(x)+{{}^{\mu}\widetilde{g}_{\lambda}}(x),$ (4.4)
@@
[…]
where, for every $(x,u)\in\mathbb{R}^{\widetilde{N}}\times\mathbb{R}^{\widetilde{N}}$,
${F(x,u)=}\widetilde{f}(x)+\widetilde{g}_{\lambda}(u)+\frac{\mu}{2}\|x-u\|^{2},$ (4.6)
and
@@
is Lipschitz-differentiable, with constant $\widetilde{\beta}+\mu$. The sequence ${(x_{k})}_{k\in\mathbb{N}}$ generated by algorithm (4.5) is then ensured to converge to the solution $x^{\ddagger}$ to (4.4), if $0<\tau<2{(\widetilde{\beta}+\mu)}^{-1}$.
In practice, as the proximity operator of $\widetilde{g}_{\frac{\lambda}{\mu}}$ often does not have a closed form, the computational complexity of this algorithm is too high. However, the solution of the proximity operator can be computed using iterative schemes such as the FB or the dual-FB algorithms, as described in Sections 2 and Sec:SD-PnP. Using a bi-level framework similar to the one proposed by [22], we show that one can use inexact proximity operator computation powered by iterative proximal algorithms with warm restarts to build an algorithm that converges to a solution to (4.4). In this context, we then focus on the following algorithm
[…]
where, for every $k\in\mathbb{N}$, $\widetilde{G}_{L,\lambda\mu^{-1}}\colon\mathbb{R}^{\widetilde{N}}\times\mathbb{R}^{\widetilde{N}}\to\mathbb{R}^{\widetilde{N}}$ is an operator that aims to approximate $\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x_{k+1})$ by computing $L$ sub-iterations of some iterative algorithm. The first input corresponds to the output $x_{k+1}$ of the gradient descent step, i.e. the point at which the proximity step should be computed. The second input corresponds to the output of $\widetilde{G}_{L,\lambda\mu^{-1}}$ from the previous iteration k1k-1, which is aimed to be used for warm-restart, i.e. initializing the sub-iterations to compute the approximation of $\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x_{k+1})$. Unlike in previous sections, in this section, this operator is not necessarily based on FB or dual-FB iterations. However, we will assume that it satisfies some sufficient decrease property (see Theorem 4.3 Condition (i)).
@@
$\text{find }x^{\ddagger}=\underset{x\in\mathbb{R}^{\widetilde{N}}}{\operatorname{argmin}}\;F(x,u^{\ddagger}_{x})\quad\text{such that}\quad u^{\ddagger}_{x}=\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x),$ (4.9)
where $F$ is defined in (4.6) and $\mu>0$. Hence, the sequence ${(x_{k})}_{k\in\mathbb{N}}$ generated by (4.5) is ensured to converge to a solution to (4.9) if $0<\tau<2{(\widetilde{\beta}+\mu)}^{-1}$.
@@
    $\phi\in\mathbb{R}^{*}\mapsto p(\phi)=(8\alpha_{L}^{2}(1-2\alpha_{L}^{2}))\phi^{2}+2(\widetilde{\beta}+\mu)(1-2\alpha_{L}^{2})\phi-\mu^{2}.$ (4.11)
@@
Proof. Let $k\in\mathbb{N}$, and let us define, for every $x\in\mathbb{R}^{\widetilde{N}}$, $h(x)=F(x,u^{\ddagger}_{x})$, where $F$ and $u^{\ddagger}_{x}\in\mathbb{R}^{\widetilde{N}}$ are defined in (4.6) and (4.9), respectively. According to (4.7), since $\nabla_{x}F(\cdot,u^{\ddagger}_{x})$ is $(\widetilde{\beta}+\mu)$-Lipschitz, then $h$ is Lipschitz-differentiable, with constant $\widetilde{\beta}+\mu>0$. Thus we can apply the descent lemma from [6, Prop. A.24] to obtain
@@
$\|u_{k+1}-u^{\ddagger}_{x_{k+1}}\|\leq\alpha_{L}\|u_{k}-u^{\ddagger}_{x_{k}}\|+\alpha_{L}\tau\|\nabla_{x}F(x_{k},u_{k})\|.$
Then, by squaring the above inequality, and by applying the Jensen’s inequality we obtain
@@
Hence, by summing (4.20) over $k\in\{0,\ldots,K-1\}$, for $K>0$, we obtain $\sum_{k=0}^{K}\frac{\tau}{2}\|\nabla h(x_{k})\|^{2}\leq\mathcal{L}_{0}-\mathcal{L}_{K}<+\infty$. Thus ${(\|\nabla h(x_{k})\|^{2})}_{k\in\mathbb{N}}$ converges to 0.
Since $h$ is convex, coercive (see [4, Cor. 11.16 & 11.17]), and has a continuous gradient, the sequence $(x_{k})_{k\in\mathbb{N}}$ is bounded and satisfies $\nabla h(x_{k})\to 0$. Therefore, all its cluster points are minimizers of $h$, and $(x_{k})$ converges to a solution to (4.9).
@@
    Condition (i) in Theorem 4.3 means that the operator $\widetilde{G}_{L,\lambda\mu^{-1}}$ in (4.8) must be chosen to satisfy some sufficient approximation condition in the sense that, for every $k\in\mathbb{N}$, $u_{k+1}=\widetilde{G}_{L,\lambda\mu^{-1}}(x_{k+1},u_{k})$ should be closer to $\operatorname{prox}_{\widetilde{g}_{\frac{\lambda}{\mu}}}(x_{k+1})$ than $u_{k}$ (i.e., the initialization of that operator).
    In practice, if we choose a monotone algorithm such as FB or dual-FB, this condition can be satisfied for an $L$ sufficiently large (due to decreasing properties on the associated objective functions). However, it can be difficult to determine the value of $L$ ensuring $\alpha_{L}<1/\sqrt{2}$.
@@
    For $(x,u)\in(\mathbb{R}^{N})^{2}$, we define $\bar{G}_{\lambda}(x,u)=\frac{1}{2}\|x-u\|^{2}+\widetilde{g}_{\lambda}(u)$. We consider the first $L$ iterates generated by an algorithm designed to estimate $u^{\ddagger}_{x}=\operatorname{prox}_{\widetilde{g}_{\lambda}}(x)=\underset{u}{\operatorname{argmin}}\;\bar{G}_{\lambda}(x,u)$, that is initialized using a warm-starting strategy, i.e., with $u_{\ell}$. Assume that the $L$-th estimate $u_{\ell+L}$ satisfies
@@
    where $C>0$ is a constant that does not depend on $x$. Then, since $\widetilde{g}_{\lambda}$ is assumed to be convex, $\bar{G}_{\lambda}(x,\cdot)$ is $1$-strongly convex [4, Prop. 10.8], and we have
    $\|u_{\ell+L}-u^{\ddagger}_{x}\|_{2}^{2}\leq\frac{2C}{L}\|u_{\ell}-u^{\ddagger}_{x}\|^{2}.$
    In this case, $L$ should be chosen larger than $4C$ to satisfy assumption (4.10). Note that this is typically the case when $g$ is the SD formulation and we use FB as an inner-solver [5].
@@
    When $\alpha_{L}$ is unknown, the root of the polynomial (4.11) cannot be computed. In this case, one can choose $\phi=1$ in the Lyapunov function (4.19), and $0<\tau<\min\Big{\{}\frac{2(1-2\alpha_{L}^{2})}{\mu^{2}},\frac{1}{\widetilde{\beta}+\mu+4\alpha_{L}^{2}}\Big{\}}$}. Since $0<\alpha_{L}^{2}<1/2$, if $\tau$ is chosen to satisfy $0<\tau<\min\big{\{}2/\mu^{2},1/(\widetilde{\beta}+\mu)\big{\}}$}, then ${(x_{k})}_{k\in\mathbb{N}}$ generated by algorithm (4.8) converges to a solution to (4.9).
[…]